If consumption is $25,0 when income is $26,000, , and consumption increases to $25,900 when income increases to $28,000 , the marginal propensity to consume is:
Answer #1

Answer: 0.45

The marginal propensity to consume is: 0.45

The marginal propensity to consume(MPC) is a metric that shows the increase in consumption for the increase in disposable income. So MPC is the ratio between the 'change in consumption' and the 'change in income'.

In the question;

Initial income (Y1) =  $26,000 , Initial consumption (C1) =  $25,000

Increased income(Y2) = $28,000 , Increased consumption(C2) = $25,900

Change in income (\DeltaY) = Y2 - Y1 = $28,000 -  $26,000 = $2,000

Change in consumption (\DeltaC) = C2 - C1 = $25,900 - $25,000 = $900

Marginal Propensity to Consume(MPC) = $900 / $2,000 = 0.45

MPC = 0.45

______________________________________________________